# Modeling and Optimizing User Preferences in AI Copilots: A Comprehensive Survey and Taxonomy

### v1

@@ 4 Preference-based Response Generation
Considering user preferences in an AI copilot as a system-level attribute that must be addressed across different stages. Fig. 1 shows the conceptual design of an AI copilot from a preference-aware perspective. It briefly illustrates the journey of user preferences—from input channels, through the development of preference representations, to their use in response generation and, finally, feedback-driven optimization. In the subsequent sections, we review related studies for each part of this preference-aware design in detail.
@@ 4.1 Preference Detection Sources
Table 2: Taxonomy of Preference Detection Sources: Mechanisms, Use Cases, and Limitations
@@ 4.1.1 Human-Driven Feedback Channels
Interactive and hybrid preference elicitation strategies aim to balance explicit clarity and implicit convenience, reducing user effort while maintaining high-quality feedback signals. One promising interactive method [43] involves self-exploring language models, which autonomously generate diverse output variations to probe user preferences. The effectiveness of such models has been demonstrated through reinforcement learning loops that iteratively refine responses based on real-time user reactions. This approach substantially reduces user effort while improving alignment accuracy. Furthermore, optimization-driven elicitation strategies utilizing Bayesian optimization have been proposed to systematically minimize user interactions. Specifically, LLM-based acquisition functions [44] have been introduced to strategically select the most informative queries, significantly decreasing interaction requirements without sacrificing the quality of preference inference. Complementary approaches have employed probabilistic reasoning [46] to dynamically adapt elicitation strategies based on uncertainty in user feedback, enabling effective handling of ambiguous interactions. In parallel, a Bayesian preference elicitation framework [45] has been developed to rigorously integrate prior knowledge with real-time user responses, enhancing the statistical validity and reliability of interactive preference modeling in conversational settings. Collectively, these interactive and hybrid elicitation methods represent a significant advancement in preference detection, optimizing user involvement and improving feedback clarity in personalized response generation systems.
@@ 4.1.2 LLM-Generated Feedback Channels
Complementing human-generated signals, recent advances increasingly leverage autonomous large language models (LLMs) as intrinsic sources of preference detection. Structured persona and profile modeling constitute a primary approach, wherein user contexts are systematically synthesized into structured intermediate representations prior to response generation. For example, a guided profile generation techniques is introduced [48], demonstrating a notable improvement in preference alignment accuracy compared to direct raw-context utilization, highlighting the efficacy of structured user profiling for personalized dialogue response generation. Optimization-driven methodologies offer another dimension of LLM-generated feedback, explicitly aligning model outputs with user preferences through targeted minimization of divergence metrics. For instance, an f-divergence minimization technique [49] is applied to finely control the alignment between user-preferred and model-generated responses, effectively balancing consistency and diversity in personalized outputs.
@@ 4.2.3 Post-Interaction Feedback and Preference Refinement
Table 3: Summary of Preference Detection Techniques Across Conversation Phases.
@@ 4.3 Personalized Response Generation
Table 4: Comparison of personalized response generation approaches organized by personalization level, mechanism, focus, advantages, and limitations.
@@ 4.4 Feedback-Driven Preference Optimization
In contrast to the approaches discussed in the previous section—which personalize responses by conditioning on explicit user preferences—another line of research focuses on post-training, preference-based alignment methods. These methods aim to align model behavior with human preferences by optimizing the model based on feedback over its generated outputs. A widely adopted approach in this category is Reinforcement Learning from Human Feedback (RLHF) [90], which involves training a reward model from human-labeled preferences and then fine-tuning the language model through reinforcement learning. In contrast, Direct Preference Optimization (DPO) [91] simplifies this process by eliminating the need for both a reward model and reinforcement learning. Instead, it directly optimizes a preference-based objective over response pairs using a supervised learning framework. In the following subsections, we review both RLHF and DPO in detail, highlighting their design choices, challenges, and recent advancements.
@@ 4.4.1 RLHF
Table 5: Summary of Key Focus Areas in the RLHF Pipeline. RLAIF stands for Reinforcement Learning from AI Feedback.
@@
A central pillar of the RLHF pipeline is reward modeling, which serves as the mechanism for translating human preferences into a trainable signal that guides policy optimization. This approach was first formalized through pairwise preference learning, in which reward models were trained using binary comparisons between model outputs, as shown in Deep RL from Human Preferences (2017) [98]. Subsequent improvements extended this method to domains such as language generation [92] and summarization [93], where human preferences were systematically incorporated to better capture quality and alignment. As the field has matured, scalability and fidelity have become pressing concerns. To address computational limitations, ensemble-based reward modeling techniques were developed to enhance performance without excessive overhead [99]. To improve feedback quality, Attention-Based Credit Assignment was proposed as a mechanism for generating dense reward signals from sparse human input [100]. In parallel, the Adaptive Preference Scaling framework introduced a dynamic way to modulate human preferences in multi-objective settings, helping models align with more nuanced value signals [101]. Conceptual challenges—such as reward hacking and misgeneralization—have also been systematically outlined in foundational critiques of RLHF reward modeling [111]. Together, these efforts reflect a clear evolution from handcrafted reward structures to scalable, robust, and statistically principled models capable of capturing complex, subjective human values—firmly establishing reward modeling as a cornerstone of alignment in RLHF.
@@ 4.4.2 DPO
Several algorithmic extensions to the original DPO formulation have been proposed to address practical limitations such as overconfidence, verbosity, prior dependency, and optimization sharpness. These are summarized in Table 6, which categorizes variants by their scoring and regularization formulations. To mitigate reward overoptimization and encourage more calibrated updates, Contrastive Preference Optimization (CPO) [112] adds a log-likelihood penalty term to the loss, discouraging the model from assigning excessively high probabilities to preferred responses. Odds-Ratio Preference Optimization (ORPO) [113], in contrast, replaces the standard log-likelihood difference with a log-odds scoring function derived from a reference model, thereby anchoring preferences relative to base policy probabilities and removing the need for explicit KL penalties. Simplex Preference Optimization (SimPO) [114] targets verbosity by scaling log-probabilities inversely with response length, effectively normalizing the preference score to favor concise outputs. Information-Ratio Preference Optimization (IRPO) [115] integrates auxiliary compression and coverage terms into both scoring and regularization components, enhancing robustness to preference noise and improving generalization under limited supervision. $\beta$-DPO [116] revisits the temperature parameter in the scoring function, replacing the fixed $\beta$ with either learnable or input-dependent temperature values to dynamically adjust alignment sensitivity across prompts. Most recently, Direct Preference Optimization with Criterion (DPOC) [63] addresses preference misranking by adding penalty terms based on criterion responses, ensuring that the preferred output meaningfully surpasses both the rejected and an intermediate-quality reference. To support controllable alignment across multiple preference dimensions—such as helpfulness, honesty, and harmlessness—multi-objective frameworks have been explored. One such approach is MODPO [117], which generalizes DPO to accommodate vectorized preference signals, enabling simultaneous optimization over multiple alignment axes without collapsing them into a single scalar reward.
@@ 5 CONCLUSION
This survey examined the role of preference optimization in the design of AI copilots—AI-powered, context-aware systems that support users in real-time, high-context tasks. We began by establishing a unified definition of AI copilots that consolidates diverse interpretations across domains, providing a conceptual basis for consistent study and development. Building on this, we offered a phase-based view of preference optimization, structured around how user preferences are identified, interpreted, and used to drive personalized interaction throughout the lifecycle of user engagement. By synthesizing methods for detecting preferences, generating contextually appropriate responses, and refining behavior through feedback, this work provides a coherent framework for understanding and advancing personalization in AI copilots. Beyond conceptual synthesis, the survey offers practical guidance for system designers by mapping available preference sources, outlining how they can be integrated, and analyzing technique selection across each stage of a preference-aware AI copilot.
As a future direction, this perspective could be extended by systematically exploring implementation pathways and cost trade-offs. Each stage involves choices—such as platform selection, model customization, or data infrastructure—that carry implications for scalability, latency, and operational cost. Studying these dimensions would help with more actionable strategies for building effective and efficient preference-aware AI copilots.

### v2

@@ 4 Preference-based Response Generation
The ability to model and respond to user preferences should be considered a system-level attribute of an AI copilot, requiring attention across multiple stages of its design and deployment. Fig. 1 shows the conceptual design of an AI copilot from a preference-aware perspective. It briefly illustrates the journey of user preferences—from input channels, through the development of preference representations, to their use in response generation and, finally, feedback-driven optimization. In the subsequent sections, we review related studies for each part of this preference-aware design in detail.
@@ 4.1 Preference Detection Sources
Table 2: Comparison of preference detection sources based on mechanisms, applications, and limitations. LLM stands for Large Language Model.
@@ 4.1.1 Human-Driven Feedback Channels
Interactive and hybrid preference elicitation strategies aim to balance explicit clarity and implicit convenience, reducing user effort while maintaining high-quality feedback signals. One promising interactive method [41] involves self-exploring language models, which autonomously generate diverse output variations to probe user preferences. The effectiveness of such models has been demonstrated through reinforcement learning loops that iteratively refine responses based on real-time user reactions. This approach substantially reduces user effort while improving alignment accuracy. Furthermore, optimization-driven elicitation strategies utilizing Bayesian optimization have been proposed to systematically minimize user interactions. Specifically, LLM-based acquisition functions [42] have been introduced to strategically select the most informative queries, significantly decreasing interaction requirements without sacrificing the quality of preference inference. Complementary approaches have employed probabilistic reasoning [44] to dynamically adapt elicitation strategies based on uncertainty in user feedback, enabling effective handling of ambiguous interactions. In parallel, a Bayesian preference elicitation framework [43] has been developed to rigorously integrate prior knowledge with real-time user responses, enhancing the statistical validity and reliability of interactive preference modeling in conversational settings. Collectively, these interactive and hybrid elicitation methods represent a significant advancement in preference detection, optimizing user involvement, and improving feedback clarity in personalized response generation systems.
@@ 4.1.2 LLM-Generated Feedback Channels
Complementing human-generated signals, recent advances increasingly leverage autonomous large language models (LLMs) as intrinsic sources of preference detection. Structured persona and profile modeling constitute a primary approach, wherein user contexts are systematically synthesized into structured intermediate representations prior to response generation. For example, a guided profile generation technique is introduced [46], demonstrating a notable improvement in preference alignment accuracy compared to direct raw-context utilization, highlighting the efficacy of structured user profiling for personalized dialogue response generation. Optimization-driven methodologies offer another dimension of LLM-generated feedback, explicitly aligning model outputs with user preferences through targeted minimization of divergence metrics. For instance, an f-divergence minimization technique [47] is applied to finely control the alignment between user-preferred and model-generated responses, effectively balancing consistency and diversity in personalized outputs.
@@ 4.2.3 Post-Interaction Feedback and Preference Refinement
Table 3: Comparison of preference detection techniques across conversation phases.
@@ 4.3 Personalized Response Generation
Table 4: Comparison of personalized response generation approaches.
@@ 4.4 Feedback-Driven Preference Optimization
In contrast to the approaches discussed in the previous section, which personalize responses by conditioning on explicit user preferences, another line of research focuses on post-training, preference-based alignment methods. These methods aim to align model behavior with human preferences by optimizing the model based on feedback over its generated outputs. A widely adopted approach in this category is Reinforcement Learning from Human Feedback (RLHF) [88], which involves training a reward model from human-labeled preferences and then fine-tuning the language model through reinforcement learning. In contrast, Direct Preference Optimization (DPO) [89] simplifies this process by eliminating the need for both a reward model and reinforcement learning. Instead, it directly optimizes a preference-based objective over response pairs using a supervised learning framework. In the following subsections, we review both RLHF and DPO in detail, highlighting their design choices, challenges, and recent advancements.
@@ 4.4.1 RLHF
Table 5: Comparison of preference alignment techniques focusing on goals, key approaches, and limitations.
@@
A central pillar of the RLHF pipeline is reward modeling, which serves as the mechanism for translating human preferences into a trainable signal that guides policy optimization. This approach was first formalized through pairwise preference learning, in which reward models were trained using binary comparisons between model outputs, as shown in Deep RL from Human Preferences [96]. Subsequent improvements extended this method to domains such as language generation [90] and summarization [91], where human preferences were systematically incorporated to better capture quality and alignment. As the field has matured, scalability and fidelity have become pressing concerns. To address computational limitations, ensemble-based reward modeling techniques were developed to enhance performance without excessive overhead [97]. To improve feedback quality, Attention-Based Credit Assignment was proposed as a mechanism for generating dense reward signals from sparse human input [98]. In parallel, the Adaptive Preference Scaling framework introduced a dynamic way to modulate human preferences in multi-objective settings, helping models align with more nuanced value signals [99]. Conceptual challenges, such as reward hacking and misgeneralization, have also been systematically outlined in foundational critiques of RLHF reward modeling [109]. Together, these efforts reflect a clear evolution from handcrafted reward structures to scalable, robust, and statistically principled models capable of capturing complex, subjective human values, thereby firmly establishing reward modeling as a cornerstone of alignment in RLHF.
@@ 4.4.2 DPO
Several algorithmic extensions to the original DPO formulation have been proposed to address practical limitations such as overconfidence, verbosity, prior dependency, and optimization sharpness. These are summarized in Table 6, which categorizes variants by their scoring and regularization formulations. To mitigate reward overoptimization and encourage more calibrated updates, Contrastive Preference Optimization (CPO) [110] adds a log-likelihood penalty term to the loss, discouraging the model from assigning excessively high probabilities to preferred responses. Odds-Ratio Preference Optimization (ORPO) [111], in contrast, replaces the standard log-likelihood difference with a log-odds scoring function derived from a reference model, thereby anchoring preferences relative to base policy probabilities and removing the need for explicit KL penalties. Simplex Preference Optimization (SimPO) [112] targets verbosity by scaling log-probabilities inversely with response length, effectively normalizing the preference score to favor concise outputs. Information-Ratio Preference Optimization (IRPO) [113] integrates auxiliary compression and coverage terms into both scoring and regularization components, enhancing robustness to preference noise and improving generalization under limited supervision. $\beta$-DPO [114] revisits the temperature parameter in the scoring function, replacing the fixed $\beta$ with either learnable or input-dependent temperature values to dynamically adjust alignment sensitivity across prompts. Most recently, Direct Preference Optimization with Criterion (DPOC) [115] addresses preference misranking by adding penalty terms based on criterion responses, ensuring that the preferred output meaningfully surpasses both the rejected and an intermediate-quality reference. To support controllable alignment across multiple preference dimensions, such as helpfulness, honesty, and harmlessness, multi-objective frameworks have been explored. One such approach is MODPO [116], which generalizes DPO to accommodate vectorized preference signals, enabling simultaneous optimization over multiple alignment axes without collapsing them into a single scalar reward.
@@ 5 CONCLUSION
In this survey, we examined the emerging landscape of preference optimization in AI-powered Cognitive Assistants and AI copilots—a domain of increasing relevance as such systems become more deeply integrated into human decision-making workflows. By offering a unified, literature-informed definition of AI copilots, we provide a conceptual foundation that bridges diverse interpretations across disciplines. Our review systematically explored the lifecycle of preference handling, from identifying relevant sources and detection techniques to mechanisms for personalized response generation and feedback-driven refinement. Collectively, these insights highlight both the progress made and the challenges that remain in designing truly adaptive and human-aligned AI assistants. Future work will benefit from continued focus on scalable, interpretable, and ethically grounded preference optimization methods that ensure alignment with user intent and values across dynamic contexts.